# Peaceful Anarcho-Accelerationism: Decentralized Full Automation for a Society of Universal Care

### v3

@@ 3 Historical and Methodological Foundations
Anarcho-accelerationism does not emerge from a vacuum. It is grounded in a 200-year intellectual tradition of anarchist political philosophy, enriched by the systematic historiography of scholars who have mapped this tradition with increasing rigor. Our methodology draws on six major historians of anarchism, namely Eltzbacher, Nettlau, Woodcock, Joll, Guérin, and Marshall, and updates their analytical categories for the technological conditions of the 21st century.
3.1 The Anarchist Historiographic Tradition
@@
Paul Eltzbacher (Eltzbacher, 1908) produced the first systematic comparative analysis of anarchist thought, examining seven thinkers, namely Godwin, Proudhon, Stirner, Bakunin, Kropotkin, Tucker, and Tolstoy, through a rigorous analytical framework that identified their positions on law, the state, and property. Eltzbacher’s method of systematic comparison across shared analytical categories remains foundational to our approach. Max Nettlau (Nettlau, 1996), known as the “Herodotus of Anarchy,” produced the most comprehensive historiographic corpus in anarchist studies, documenting movements across continents and centuries in his multi-volume Geschichte der Anarchie (1925–1931). Nettlau’s contribution is methodological: he demonstrated that anarchism is not a singular doctrine but a family of movements unified by shared opposition to domination and commitment to voluntary association.
George Woodcock (Woodcock, 1962) provided the standard English-language history, periodizing anarchist thought from Godwin’s Enquiry Concerning Political Justice (Godwin, 1793) through the classical period of Proudhon (Proudhon, 1840), Bakunin, and Kropotkin, to the Spanish Revolution and beyond. James Joll (Joll, 1964) situated anarchism within the broader context of European political movements, analyzing it as a recurring response to industrialization and state centralization. Daniel Guérin (Guérin, 1970) synthesized anarchist and Marxist insights, emphasizing anarchism as a practical movement oriented toward workers’ self-management rather than a purely philosophical current. Peter Marshall (Marshall, 1992) produced the most comprehensive single-volume history, tracing libertarian ideas from ancient Taoism and Greek Stoicism through the Enlightenment to contemporary movements, demonstrating the planetary scope and historical depth of the anarchist impulse.
3.2 Seven Categories of Anarchism Updated for Peaceful Accelerationism
Drawing on Eltzbacher’s comparative method and enriched by the subsequent historiographic tradition, we identify seven analytical categories through which anarchist thought can be systematically examined. Table˜3 presents these categories as they appear in classical anarchism and as we update them for peaceful anarcho-accelerationism.
@@ 3.2 Seven Categories of Anarchism Updated for Peaceful Accelerationism
Table 3: Seven analytical categories of anarchism, updated from classical anarchism to peaceful anarcho-accelerationism.
@@
Social classes Bourgeoisie vs. proletariat; ruling class vs. working class Cloud lords vs. cloud serfs; platform owners vs. platform-dependent precariat; data extractors vs. data producers
@@
The update is not a rupture with classical anarchism but its logical extension to new conditions of domination. Where Proudhon confronted industrial capital, we confront platform capital. Where Bakunin organized workers against the factory owner, we organize communities against the algorithm. Where Kropotkin documented mutual aid in nature, we document mutual aid in open-source software. The principle remains constant: self-organization without domination. The terrain has shifted from the factory floor to the digital commons.
3.3 Ten Ethical Principles of Peaceful Anarcho-Accelerationism
Drawing on Kropotkin’s naturalistic ethics (Kropotkin, 1924), Tolstoy’s principled nonviolence (Tolstoy, 1894), Gandhi’s satyagraha (Gandhi, 1961), and the care ethics tradition (Tronto, 1993), we articulate ten ethical principles that govern the peaceful anarcho-accelerationist project. These principles are not abstract ideals but design constraints for the Liberation Stack, and every technology, protocol, and institution proposed in this paper can be evaluated against them.
@@ 3.3 Ten Ethical Principles of Peaceful Anarcho-Accelerationism
The first is non-domination: no person, institution, or algorithm shall exercise coercive power over another, and technology must distribute power, never concentrate it. The second is universal care: every system, institution, and technology shall be evaluated by its capacity to sustain life, relationships, and communities, since care is the organizing principle, not profit. The third is nonviolence: social transformation shall be achieved through peaceful means, including technological commons-building, education, mutual aid, prefigurative politics, and democratic persuasion, following Gandhi (Gandhi, 1961) and Sharp (Sharp, 1973). The fourth is commons governance: essential resources such as energy, food, knowledge, communication, code, and productive infrastructure shall be governed as commons according to Ostrom’s principles (Ostrom, 1990), not enclosed as property. The fifth is ecological integrity: all productive activity shall operate within planetary boundaries (Raworth, 2017), since sustainability is a first-order constraint, not an externality.
The sixth is radical transparency: algorithms that affect human lives shall be auditable, data collection shall be consensual, and governance shall be open, because opacity is a tool of domination. The seventh is solidarity across difference: following intersectional analysis (Crenshaw, 1989; Collins, 2019), liberation must be universal, since no one is free until everyone is free, and UDR eliminates the material basis of all axes of oppression. The eighth is intellectual commitment: following Chomsky’s principle of the responsibility of intellectuals (Chomsky, 1967), those with knowledge and privilege have a heightened obligation to speak truth to power and to build alternatives. The ninth is freedom of conscience: every person’s spiritual, philosophical, and existential commitments shall be respected, since the framework serves human flourishing in all its dimensions, including the spiritual (Lumbreras and Garrido-Merchán, 2024). The tenth is prefigurative consistency: the means must embody the ends, systems designed for democracy must be governed democratically, and tools for liberation must not reproduce domination, as Ward wrote, “anarchy in action” (Ward, 1973).
4 The Bridge-Building Thesis
@@
The most distinctive claim of anarcho-accelerationism is that it does not seek to defeat other traditions but to find common ground with all of them. Table˜4 summarizes the convergence. We elaborate each bridge in turn.
@@
Liberalism’s greatest insight is that individual freedom matters. Anarcho-accelerationism shares this commitment without reservation. Where we diverge is on the conditions for genuine autonomy. A person who depends on a monopolistic employer for income, a monopolistic landlord for shelter, and a monopolistic platform for communication is not free in any meaningful sense (Pettit, 1997). Genuine autonomy requires material independence, which in turn requires commons-based infrastructure: community-owned solar microgrids that eliminate dependence on centralized utilities (Farrell, 2016), self-hosted federated services (Mastodon, Matrix, Nextcloud) that eliminate dependence on corporate platforms (Zulli et al., 2020), and platform cooperatives (Fairbnb, Stocksy, Up & Go) that eliminate dependence on extractive intermediaries (Scholz, 2016). The liberal who genuinely cares about individual freedom should welcome the anarcho-accelerationist program: more autonomy, not less, built on shared infrastructure rather than corporate dependency.
@@
Socialism’s greatest insight is that material inequality undermines both freedom and democracy. Anarcho-accelerationism shares this commitment. Where we diverge is on the mechanism. State ownership historically concentrates power in a bureaucratic class that reproduces inequality in new forms (Scott, 1998). Commons-based ownership, where resources are governed by their users according to Ostrom’s principles, distributes power more effectively than either state or market. The Mondragon cooperative corporation, with approximately 70,000 worker-owners and over € 11 billion in annual revenue, demonstrates that democratic enterprise is economically viable at industrial scale (Whyte and Whyte, 1991). The Emilia-Romagna region of Italy, where cooperative enterprises account for 30% of GDP, consistently ranks among the highest in quality of life in Europe (Restakis, 2010). The socialist who genuinely cares about equality should welcome the anarcho-accelerationist program: deeper equality, achieved through distributed ownership rather than centralized redistribution.
Environmentalism’s greatest insight is that economic activity must respect biophysical limits. Anarcho-accelerationism integrates this as a first-order constraint. But we go further: the ecological crisis is structurally inseparable from the concentration of power. Just 100 companies are responsible for 71% of global industrial greenhouse gas emissions (Griffin, 2017). Fossil fuels are inherently centralizing (extraction requires massive capital, refining requires specialized facilities, distribution requires controlled networks), while renewable energy is inherently decentralizing (sunlight and wind are distributed, modular, and scalable) (Rifkin, 2011). Local food systems can significantly reduce the transport component of food-related greenhouse gas emissions, though production-phase emissions (accounting for 83% of the total) typically dominate (Weber and Matthews, 2008), and agroecological methods sequester carbon and restore biodiversity (Altieri and Nicholls, 2012). Figure˜3 illustrates the dramatic decline in solar energy costs over the past decade, demonstrating that the economics of decentralized renewable energy have already surpassed those of centralized fossil fuels. The environmentalist who genuinely cares about the planet should welcome the anarcho-accelerationist program: systemic sustainability, achieved through decentralized infrastructure rather than top-down regulation of fundamentally unsustainable systems.
@@
Figure 3: Global weighted-average levelized cost of energy (LCOE) for utility-scale solar photovoltaics compared to natural gas, 2010–2022. Data from IRENA (IRENA, 2023). Solar PV costs have declined by approximately 89% in twelve years, crossing below fossil fuel costs around 2019. This cost inversion makes decentralized community-owned energy not only environmentally necessary but economically optimal.
@@
Feminism’s greatest insight, particularly in its materialist and care-ethics strands, is that the work of sustaining life (caring, teaching, healing, nurturing, feeding, cleaning) is the foundation of all economic activity, yet it is systematically devalued and rendered invisible (Federici, 2012; Tronto, 1993). Anarcho-accelerationism places care at the center of its value system, not as a rhetorical gesture but as a design principle. Systems should be evaluated by their care capacity: do they help people take care of each other, or do they fragment communities? Automation should liberate time for care, not replace care with machines. The labor of maintaining systems (emotional labor, maintenance work, moderation) should be visible, valued, and equitably distributed (Star, 1999). The feminist who genuinely cares about dismantling patriarchal structures should welcome the anarcho-accelerationist program: a sociotechnical order where care is not a burden borne disproportionately by women but the organizing principle of collective life.
The hacker ethic, articulated by Levy (Levy, 2010) and formalized by Stallman’s four freedoms (Stallman, 2002), holds that information should be free, that access to computers should be unlimited, and that decentralization promotes liberty. Anarcho-accelerationism is the political extension of this ethic beyond software: if code should be free, why not seeds? (Open Source Seed Initiative (Kloppenburg, 2014)). If protocols should be open, why not energy grids? (Farrell, 2016). If platforms should be interoperable, why not governance? (Decidim, Loomio, Pol.is as open-source democracy tools (Barandiaran et al., 2024)). The hacker who genuinely believes in the four freedoms should welcome the anarcho-accelerationist program: the universalization of open-source principles to all domains of collective life.
The concept of the common good has deep roots in Catholic social teaching (Pope Francis, 2015), Islamic economics (waqf, zakat), Buddhist economics (Schumacher, 1973), and indigenous cosmovisions (buen vivir, sumak kawsay) (Gudynas, 2011). These traditions share a conviction that the goods of the Earth are destined for all, that human dignity is non-negotiable, and that community is prior to the individual. Anarcho-accelerationism speaks this language without requiring religious commitment: commons governance operationalizes the “universal destination of goods,” care ethics operationalizes “love thy neighbor,” and ecological limits operationalize “stewardship of creation.” This bridge is particularly significant in Latin America, Southern Europe, and the Global South, where faith-based communities are often the strongest existing networks of mutual aid (Deneulin and Bano, 2009). A framework that can speak to both the hacker and the parish, the cooperative and the congregation, the fablab and the food bank, has a much wider reach than one that demands ideological purity.
Crucially, anarcho-accelerationism respects the freedom of faith not merely as a political concession but on principled epistemological and ontological grounds. As argued by Lumbreras and Garrido-Merchán (2024), the ontological scope of theology, the nature of God, the soul, grace, and mystical union, categorically escapes the epistemological reach of artificial general intelligence (AGI). Through the frameworks of Saint Teresa of Avila’s Interior Castle (Las Moradas) and Saint Augustine’s Confessions, Lumbreras and Garrido-Merchán demonstrate that human interiority, the layered structure of the soul oriented toward the divine, is irreducible to computational processes. Saint Teresa maps seven successive moradas (dwelling places) of the soul, each representing a deeper level of self-knowledge and encounter with God, culminating in mystical union. Saint Augustine’s principle that “truth dwells in the inner man” (in interiore homine habitat veritas) grounds genuine knowledge in a first-person subjective perspective that no information-processing system can replicate. AGI can simulate theological discourse, generate text about spiritual experience, and model the behavioral correlates of religious practice, but it cannot undergo the interior journey that Teresa describes, because that journey requires phenomenal consciousness, which, as we have argued (Section˜2), is independent of computational intelligence (Garrido-Merchán and Lumbreras, 2022). The theological domain, precisely because it concerns realities accessible only through interiority and faith, lies beyond the epistemological horizon of any computational system, no matter how powerful. This means that full automation and AGI do not threaten authentic religious life; they liberate time for it. A person freed from instrumental labor by the robotic commons has more time for prayer, contemplation, community worship, charitable service, and the cultivation of the interior life that Teresa and Augustine describe. Anarcho-accelerationism, far from being incompatible with faith, creates the material conditions for the deepest spiritual flourishing.
@@
Table 4: Bridge-building: shared values between anarcho-accelerationism (AA) and major traditions.
Tradition Core Value Shared Principle AA Contribution
@@
Faith traditions Common good Shared dignity, mutual aid Operationalization through commons governance; freedom of faith preserved
@@
System of Domination State coercion Techno-feudalism (Varoufakis, 2023) Capitalist vassalage Surveillance capitalism (Zuboff, 2019) Social Transformation Commons-building Open-source infrastructure Peaceful transition UBI $\to$ UDR Federated governance System of Self-Management AI-assisted governance Robotic commons Community sovereignty Ecological sustainability Universal care CritiqueStrategy Based on critique of domination (Eltzbacher, Nettlau, Woodcock, Joll, Guérin, Marshall) Peaceful accelerationist commons-building strategy (Liberation Stack) Defense of self-management (AI + robotics under commons governance) The Anarcho-Accelerationist Process
[…]
Figure 5: Chronology of the three intellectual traditions that converge in peaceful anarcho-accelerationism (PAA): the anarchist tradition (from Godwin through Rojava), the accelerationist/post-scarcity tradition (from Bookchin through Srnicek and Williams), and the pacifist tradition (from Tolstoy through Gandhi, Sharp, and Chomsky). PAA (2026) synthesizes all three currents.
5 From Intersectional Liberation to Human Flourishing
The bridge-building thesis (Section˜4) demonstrates that anarcho-accelerationism finds common ground with existing traditions. In this section we go further: we show that decentralized full automation with Universal Desired Resources (UDR) achieves what no previous framework has accomplished, the simultaneous elimination of all axes of structural oppression identified by intersectional theory, the universal satisfaction of basic human needs as described by Maslow’s hierarchy, and the provision of real freedom (not merely formal freedom) through a progressive transition from the current system. This is not incremental reform; it is the material basis for universal human flourishing.
5.1 Intersectional Liberation Through UDR
Intersectionality, as articulated by Crenshaw (1989) and developed by Collins (2019), identifies that systems of oppression do not operate independently but intersect: race, gender, class, sexuality, disability, age, nationality, and other axes of identity produce compound disadvantages that cannot be understood by examining any single dimension in isolation. A Black disabled woman does not experience racism, ableism, and sexism separately; she experiences a unique, compounded form of oppression that is qualitatively different from the sum of its parts. The “matrix of domination” (Collins, 2019) describes how these intersecting systems are sustained by institutional structures, cultural norms, and, crucially, material conditions: access to employment, housing, healthcare, education, food, and social participation.
We argue that UDR, implemented through the decentralized robotic commons, constitutes the most radical intersectional intervention ever proposed, because it eliminates the material basis of all axes of oppression simultaneously. Consider each axis in turn.
@@ 5.1 Intersectional Liberation Through UDR
UDR eliminates class by definition: when every person has direct access to the goods and services they need from the robotic commons, the distinction between those who own the means of production and those who must sell their labor dissolves entirely. There is no proletariat and no bourgeoisie in a society where production is automated and commons-governed. Racial oppression, sustained by discriminatory hiring, housing segregation, unequal access to education and healthcare, and the wealth gap produced by centuries of extraction, loses its material leverage when employment, housing, food, healthcare, and education are provided by the commons independently of market participation. A person’s race ceases to determine their access to resources. The gendered division of labor, with its systematic devaluation of care work performed disproportionately by women (Federici, 2012), loses its material foundation when all instrumental labor is automated; care remains, but it is no longer an economic burden, it is a freely chosen activity in a society organized around care as its central value. When access to resources is decoupled from productive capacity, the distinction between “able” and “disabled” bodies loses its economic significance, and assistive technologies, produced on demand by the commons, become available to all who need them. Economic dependence on employers and landlords who may discriminate on the basis of sexual orientation or gender identity is eliminated when material needs are met through the commons, so that LGBTQ+ individuals no longer face the economic coercion that forces them into closeted lives or precarious employment. Both the young and the elderly, currently marginalized by their exclusion from wage labor, gain full access to the resources of the commons, valued not for their productivity but for what they are: persons deserving of care. And when material well-being is no longer tied to citizenship in a wealthy state, the economic desperation that drives forced migration is eliminated, so that people move freely because they choose to, not because survival requires it.
The key insight is structural: intersectional oppression is maintained through material scarcity administered through markets. When markets are superseded by the robotic commons under UDR, the material infrastructure of oppression is dismantled at its foundation. This does not mean that cultural prejudices disappear automatically; they do not. But it means that prejudice loses its institutional teeth: the capacity to deny resources, opportunities, and dignity on the basis of identity. Cultural transformation follows material transformation, as the historical record of post-emancipation, post-suffrage, and post-civil-rights societies demonstrates, imperfectly but unmistakably.
Table˜5 summarizes how UDR addresses each axis of intersectional oppression.

### v5

@@ 2 Theoretical Foundations: A Formal Definition of Peaceful Anarcho-Accelerationism
Anarcho-accelerationism does not emerge from a vacuum. It is grounded in a 200-year intellectual tradition of anarchist political philosophy, enriched by the systematic historiography of scholars who have mapped this tradition with increasing rigor. Our methodology draws on six major historians of anarchism and updates their analytical categories for the technological conditions of the 21st century, following the principle of totality and interdependence articulated by Corrêa (Corrêa, 2015).
@@
Paul Eltzbacher (Eltzbacher, 1908) produced the first systematic comparative analysis of anarchist thought, examining seven thinkers, namely Godwin, Proudhon, Stirner, Bakunin, Kropotkin, Tucker, and Tolstoy, through a rigorous analytical framework that identified their positions on law, the state, and property. Eltzbacher’s method of systematic comparison across shared analytical categories remains foundational to our approach. Max Nettlau (Nettlau, 1996), known as the Herodotus of Anarchy, produced the most comprehensive historiographic corpus in anarchist studies, documenting movements across continents and centuries in his multi-volume Geschichte der Anarchie. Nettlau’s contribution is methodological: he demonstrated that anarchism is not a singular doctrine but a family of movements unified by shared opposition to domination and commitment to voluntary association. George Woodcock (Woodcock, 1962) provided the standard English-language history, periodizing anarchist thought from Godwin’s Enquiry Concerning Political Justice (Godwin, 1793) through the classical period to the Spanish Revolution and beyond. James Joll (Joll, 1964) situated anarchism within the broader context of European political movements. Daniel Guérin (Guérin, 1970) synthesized anarchist and Marxist insights, emphasizing anarchism as a practical movement oriented toward workers’ self-management. Peter Marshall (Marshall, 1992) produced the most comprehensive single-volume history, tracing libertarian ideas from ancient Taoism through the Enlightenment to contemporary movements.
Peaceful anarcho-accelerationism takes specific contributions from each of Eltzbacher’s seven thinkers, as summarized in Table˜3. From Godwin (Godwin, 1793) it takes reason as the instrument of social progress: the conviction that rational inquiry, not authority, should guide human organization, and that education and rational discourse are the primary mechanisms of social change. From Proudhon (Proudhon, 1840) it takes the critique of property as monopoly and the proposal of mutualism, the blueprint for commons-based economic organization, and the principle of federation as the alternative to both state centralism and market atomization. From Stirner (Stirner, 2014) it takes the insistence that liberation must serve the concrete individual, not an abstract collective, a corrective against the authoritarian temptation in all revolutionary movements. From Bakunin (Bakunin, 1990) it takes the fundamental anarchist thesis that no state apparatus, however well-intentioned, can serve as an instrument of liberation, which grounds the insistence on stateless, federated governance. From Kropotkin (Kropotkin, 1902) it takes the scientific grounding of cooperation as a factor of evolution, the ethics of mutual aid, and the vision of a post-scarcity society achieved through the rational application of technology. From Tucker (Tucker, 1893) it takes the libertarian commitment to eliminating structural monopolies, updated for the 21st century as the monopolies of platform data, proprietary AI, centralized energy, and intellectual property. From Tolstoy (Tolstoy, 1894) it takes the absolute commitment to peaceful means, the rejection of revolutionary violence, and the conviction that the means must be consistent with the ends.
Table 3: Eltzbacher’s seven thinkers and the specific contribution of each to peaceful anarcho-accelerationism.
Thinker Period Contribution to peaceful anarcho-accelerationism
Godwin 1756–1836 Reason as instrument of social progress; education and persuasion over coercion
Proudhon 1809–1865 Mutualism; federation as alternative to state and market; commons-based economy
Stirner 1806–1856 Individual sovereignty; liberation must serve the concrete person, not abstractions
Bakunin 1814–1876 Anti-statism; no state apparatus can serve as instrument of liberation
Kropotkin 1842–1921 Mutual aid as evolutionary factor; post-scarcity through rational technology
Tucker 1854–1939 Elimination of structural monopolies (updated: data, AI, energy, IP)
Tolstoy 1828–1910 Principled nonviolence; means must be consistent with ends
Following the methodology of Corrêa (Corrêa, 2015), who insists on the principle of totality and interdependence in analyzing anarchist movements, we define peaceful anarcho-accelerationism through seven analytical categories that are not intelligible in isolation but condition and are conditioned by each other in a dialectical whole. The seven categories, compared with their classical formulations, are presented in Table˜4.
@@
Table 4: Seven analytical categories of anarchism, updated from classical anarchism to peaceful anarcho-accelerationism, following Eltzbacher’s comparative method and Corrêa’s principle of totality.
@@
Social classes Bourgeoisie vs. proletariat Cloud lords vs. cloud serfs; platform owners vs. platform-dependent precariat; data extractors vs. data producers
Corrêa’s methodology requires that these categories be understood through the principle of totality and interdependence across three registers. The first register concerns the unity of theory, practice, and history. Peaceful anarcho-accelerationism is simultaneously a theoretical framework (a formal analysis of domination and liberation under conditions of full automation), a practical program (the Liberation Stack, the UBI-to-UDR transition, the nonviolent mobilization strategy), and a historical continuation of the anarchist movement (from Godwin’s rational anarchism through the Spanish collectives to the Fediverse). No register is intelligible without the others: theory without practice is academic speculation, practice without theory is blind activism, and both without historical grounding repeat the errors of the past. The second register concerns the unity of thought and action. The intellectual analysis of domination (identifying techno-feudalism, surveillance capitalism, algorithmic management as the contemporary forms of the state-capital nexus) must be inseparable from the construction of alternatives (building community energy cooperatives, deploying federated communication platforms, establishing open-source manufacturing). In the language of Corrêa’s especifismo, the political level of theoretical analysis must be embedded within the social level of popular movements and concrete organizing. This is what Corrêa calls social insertion: the practice of embedding anarchist principles within existing social movements rather than standing apart from them in ideological purity. The third register concerns the unity of form and content. The organizational form of the movement, federated, horizontal, commons-based, must reflect its content, which is liberation, care, and non-domination. A movement that advocates decentralization cannot be organized hierarchically. A framework that promotes care cannot treat its participants instrumentally. This is the principle of prefigurative consistency (Graeber, 2013): the means must embody the ends.
Corrêa further identifies social vectors, the specific social forces that carry the potential for transformation, and critiques, the specific structures of domination against which the movement is directed. In our framework, the social vectors are the precariat (the growing class of workers in precarious, platform-mediated employment), the open-source development community (which already practices commons-based peer production at planetary scale), the cooperative movement (from Mondragon to platform cooperatives), the agroecological movement, the community energy movement, and the Fediverse community (which builds and maintains decentralized communication infrastructure). The critiques are directed against techno-feudalism (the concentration of digital productive capacity in a handful of platform corporations), surveillance capitalism (the extraction and monetization of behavioral data), algorithmic management (the replacement of human supervisors with opaque automated systems), fossil-fuel centralization (the structural dependence on concentrated energy sources that enables authoritarian control), and intellectual property enclosure (the artificial scarcification of knowledge, seeds, and code through patent and copyright regimes).
@@
The update from classical to accelerationist anarchism is not a rupture with the tradition but its logical extension to new conditions of domination. Where Proudhon confronted industrial capital, we confront platform capital. Where Bakunin organized workers against the factory owner, we organize communities against the algorithm. Where Kropotkin documented mutual aid in nature, we document mutual aid in open-source software. The principle remains constant: self-organization without domination. The terrain has shifted from the factory floor to the digital commons.
@@
Drawing on Kropotkin’s naturalistic ethics (Kropotkin, 1924), Tolstoy’s principled nonviolence (Tolstoy, 1894), Gandhi’s satyagraha (Gandhi, 1961), and the care ethics tradition (Tronto, 1993), we articulate ten ethical principles that govern the peaceful anarcho-accelerationist project. These principles are not abstract ideals but design constraints for the Liberation Stack, and every technology, protocol, and institution proposed in this paper can be evaluated against them. The first is non-domination: no person, institution, or algorithm shall exercise coercive power over another, and technology must distribute power, never concentrate it. The second is universal care: every system, institution, and technology shall be evaluated by its capacity to sustain life, relationships, and communities, since care is the organizing principle, not profit. The third is nonviolence: social transformation shall be achieved through peaceful means, including technological commons-building, education, mutual aid, prefigurative politics, and democratic persuasion, following Gandhi (Gandhi, 1961) and Sharp (Sharp, 1973). The fourth is commons governance: essential resources such as energy, food, knowledge, communication, code, and productive infrastructure shall be governed as commons according to Ostrom’s principles (Ostrom, 1990), not enclosed as property. The fifth is ecological integrity: all productive activity shall operate within planetary boundaries (Raworth, 2017), since sustainability is a first-order constraint, not an externality. The sixth is radical transparency: algorithms that affect human lives shall be auditable, data collection shall be consensual, and governance shall be open, because opacity is a tool of domination. The seventh is solidarity across difference: following intersectional analysis (Crenshaw, 1989; Collins, 2019), liberation must be universal, since no one is free until everyone is free. The eighth is intellectual commitment: following Chomsky’s principle of the responsibility of intellectuals (Chomsky, 1967), those with knowledge and privilege have a heightened obligation to speak truth to power and to build alternatives. The ninth is freedom of conscience: every person’s spiritual, philosophical, and existential commitments shall be respected, since the framework serves human flourishing in all its dimensions, including the spiritual (Lumbreras and Garrido-Merchán, 2024). The tenth is prefigurative consistency: the means must embody the ends, systems designed for democracy must be governed democratically, and tools for liberation must not reproduce domination (Ward, 1973).
Figure 3: Chronology of the three intellectual traditions that converge in peaceful anarcho-accelerationism (PAA): the anarchist tradition (from Godwin through Rojava), the accelerationist tradition (from Bookchin through Srnicek and Williams), and the pacifist tradition (from Tolstoy through Gandhi, Sharp, and Chomsky).
Anarcho-accelerationism rests on four intellectual pillars that bridge the historical and the contemporary. Kropotkin’s Mutual Aid (Kropotkin, 1902) challenged the Social Darwinist misreading of natural selection by documenting extensive cooperation across species. Contemporary evolutionary biology has vindicated this view: multilevel selection theory (Wilson, 2007), reciprocal altruism (Trivers, 1971), and cultural group selection (Henrich, 2016) all demonstrate that cooperation is a primary driver of complex social organization. In computer science, this principle is realized in commons-based peer production (Benkler, 2006). Ostrom’s institutional analysis (Ostrom, 1990) demolished Hardin’s tragedy of the commons by documenting thousands of cases where communities successfully manage shared resources without either privatization or state control. Her eight design principles provide a rigorous institutional framework that maps directly onto distributed system design (De Moor, 2010). Following Stiegler (Stiegler, 2010), we adopt the concept of the pharmakon: every technology is simultaneously remedy and poison. This is not a reason for technological pessimism; it is a reason for technological governance (Feenberg, 2002). Drawing from the anarchist tradition of prefigurative politics (Graeber, 2013), we propose that the systems we build should embody the values they aim to promote.
3 Bridge-Building and Intersectional Liberation
@@
The most distinctive claim of anarcho-accelerationism is that it does not seek to defeat other traditions but to find common ground with all of them. Table˜5 summarizes the convergence.
@@
Liberalism’s greatest insight is that individual freedom matters. Anarcho-accelerationism shares this commitment without reservation. Where we diverge is on the conditions for genuine autonomy. A person who depends on a monopolistic employer for income, a monopolistic landlord for shelter, and a monopolistic platform for communication is not free in any meaningful sense (Pettit, 1997). Genuine autonomy requires material independence, which in turn requires commons-based infrastructure: community-owned solar microgrids (Farrell, 2016), self-hosted federated services that eliminate dependence on corporate platforms (Zulli et al., 2020), and platform cooperatives that eliminate dependence on extractive intermediaries (Scholz, 2016).
@@
Socialism’s greatest insight is that material inequality undermines both freedom and democracy. Commons-based ownership, where resources are governed by their users according to Ostrom’s principles, distributes power more effectively than either state or market. The Mondragon cooperative corporation, with approximately 70,000 worker-owners and over € 11 billion in annual revenue, demonstrates that democratic enterprise is economically viable at industrial scale (Whyte and Whyte, 1991). The Emilia-Romagna region of Italy, where cooperative enterprises account for 30% of GDP, consistently ranks among the highest in quality of life in Europe (Restakis, 2010).
Environmentalism’s greatest insight is that economic activity must respect biophysical limits. The ecological crisis is structurally inseparable from the concentration of power: just 100 companies are responsible for 71% of global industrial greenhouse gas emissions (Griffin, 2017). Fossil fuels are inherently centralizing, while renewable energy is inherently decentralizing (Rifkin, 2011). Figure˜4 illustrates the dramatic decline in solar energy costs.
@@
Figure 4: Global weighted-average levelized cost of energy for utility-scale solar PV compared to natural gas, 2010–2022. Data from IRENA (IRENA, 2023). Solar costs have declined by approximately 89% in twelve years.
@@
Feminism’s greatest insight, particularly in its materialist and care-ethics strands, is that the work of sustaining life is the foundation of all economic activity, yet it is systematically devalued (Federici, 2012; Tronto, 1993). Anarcho-accelerationism places care at the center of its value system not as a rhetorical gesture but as a design principle. Automation should liberate time for care, not replace care with machines.
The hacker ethic, articulated by Levy (Levy, 2010) and formalized by Stallman’s four freedoms (Stallman, 2002), holds that information should be free, access to computers unlimited, and decentralization promotes liberty. Anarcho-accelerationism is the political extension of this ethic beyond software. If code should be free, why not seeds? (Kloppenburg, 2014). If protocols should be open, why not energy grids? (Farrell, 2016). The Fediverse, built on the ActivityPub protocol and comprising Mastodon, PeerTube, PixelFed, Lemmy, Funkwhale, and BookWyrm, represents the most complete existing implementation of the hacker ethic applied to social communication. It demonstrates that a social communication platform can operate at scale, hosting over 10 million users across thousands of independently operated servers, without advertising, without algorithmic manipulation, and without centralized ownership (Zulli et al., 2020). Each server is a sovereign community with its own rules, moderation policies, and funding model, yet all communicate through a shared open standard. The Fediverse is anarchism in practice: federated, self-governing, commons-based, and voluntary.
The concept of the common good has deep roots in Catholic social teaching (Pope Francis, 2015), Islamic economics, Buddhist economics (Schumacher, 1973), and indigenous cosmovisions (Gudynas, 2011). Anarcho-accelerationism speaks this language without requiring religious commitment. Crucially, it respects the freedom of faith not merely as a political concession but on principled epistemological and ontological grounds. As argued by Lumbreras and Garrido-Merchán (2024), the ontological scope of theology categorically escapes the epistemological reach of artificial general intelligence. Through the frameworks of Saint Teresa of Avila’s Interior Castle and Saint Augustine’s Confessions, they demonstrate that human interiority is irreducible to computational processes. Full automation and AGI do not threaten authentic religious life; they liberate time for it.
@@
Table 5: Bridge-building: shared values between anarcho-accelerationism and major traditions.
Tradition Core Value Shared Principle Contribution
@@
Faith traditions Common good Shared dignity, mutual aid Commons governance; freedom of faith preserved
@@
Intersectionality, as articulated by Crenshaw (1989) and developed by Collins (2019), identifies that systems of oppression do not operate independently but intersect: race, gender, class, sexuality, disability, age, nationality, and other axes of identity produce compound disadvantages that cannot be understood by examining any single dimension in isolation. We argue that UDR, implemented through the decentralized robotic commons, constitutes the most comprehensive intersectional intervention yet proposed, because it eliminates the material basis of all axes of oppression simultaneously. UDR eliminates class by definition: when every person has direct access to goods and services from the robotic commons, the distinction between those who own the means of production and those who must sell their labor dissolves. Racial oppression loses its material leverage when employment, housing, food, healthcare, and education are provided by the commons independently of market participation. The gendered division of labor loses its material foundation when all instrumental labor is automated. When access to resources is decoupled from productive capacity, the distinction between able and disabled bodies loses its economic significance. Economic dependence on discriminatory employers is eliminated for LGBTQ+ individuals. Both the young and the elderly gain full access to the resources of the commons. And when material well-being is no longer tied to citizenship in a wealthy state, the economic desperation that drives forced migration is eliminated. Table˜6 summarizes these interventions.